OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -9 x^3 + 8x

OpenStudy (anonymous):

help!

OpenStudy (campbell_st):

for an even function substitute -x and you should get the orginal function e.g. f(x) = x^2 + 5 f(-x) = (-x)^2 + 5 = x^2 + 5 so its even for odd functions ... f(-x) = - f(x) since your function has a highest power of 3 it may be odd so substitute x = -x and check f(-x) = -9(-x)^3 + 8(-x) when you simplify it, will you get the negative version of the original function..?

OpenStudy (campbell_st):

a function that is neither odd nor even... won't be f(x) or -f(x) when x = -x is substituted. e.g. f(x) = x^2 + 2x f(-x) = (-x)^2 + 2(-x) = x^2 - 2x which is neither odd nor even. hope this helps

OpenStudy (anonymous):

@campbell_st oh so it's neither?

OpenStudy (campbell_st):

so looking at the problem f(-x) = -9(-x)^3 + 8(-x) = -9(-x^3) - 8x = 9x^3 - 8x = -(-9x^3 + 8x) so f(-x) = - f(x) so the function is odd.

OpenStudy (anonymous):

thanku!!!:)